[epi2] Epi2 Chamber A update 1/23/13

maurice stevens maurice at stanford.edu
Wed Jan 23 14:43:50 PST 2013


We are still battle pressure issues.  Initially the system was not pumping to 15 torr with process gases flowing.  The pump was sent  out the end of last week and we pulled the pump from Chamber B.  We tried to run over the weekend but the pressure was not controlling…it would reach the 15 torr setpoint  but would then oscillate out of control.

Today Ted had Sean from AMAT back in and they have ruled out the throttle valve.  They did find that voltage coming off the pressure manometer is not stable.  Sean has ordered a new manometer and expects it to be here tomorrow.



More information about the epi2 mailing list